OpenStudy (anonymous):

Solve –2x2 +3x – 9 = 0. (solving quadratic equation with complex numbers).

OpenStudy (jamesj):

Ok, so use the quadratic formula. What is that formula?

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}\]

OpenStudy (jamesj):

right. So apply that formula with your equation.

OpenStudy (anonymous):

remember how I said last time to find out what a b and c are and plug them into the formula

OpenStudy (anonymous):

yah.

OpenStudy (anonymous):

You will get two answers for x \[\pm \] That signs means you have to do it two times. One time with + and then a second time with -

OpenStudy (anonymous):

okay , so first I factor 3x?

OpenStudy (anonymous):

Is this your equation?\[–2x^2 +3x – 9 = 0\]

OpenStudy (anonymous):

yes, srry I forgot to put ^.

OpenStudy (anonymous):

Ok you don't need to factor or do anything at all to find out what a b and c are \[–2x^2 +3x – 9 = 0\] then a b and c are \[ax^2 +bx + c = 0\]

OpenStudy (anonymous):

I'll give you a to help you out \[a=-2\]

OpenStudy (anonymous):

can you tell me what b and c are?

OpenStudy (jamesj):

Very nice Romero.

OpenStudy (anonymous):

B is 3x and C is 9?

OpenStudy (anonymous):

We just want numbers for a b and c so no x

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so\[b=3\]

OpenStudy (anonymous):

oh so I was right

OpenStudy (anonymous):

just with out the X

OpenStudy (anonymous):

also see how I also included the negative sign in a \[a=-2\] you also have to included for c

OpenStudy (anonymous):

Yeah and you forgot the negative sign for c

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

\[a=-2\] \[b=3\] \[c=-9\] correct?

OpenStudy (anonymous):

yep.

OpenStudy (jamesj):

So now, apply the formula Romero gave you above. What do you get?

OpenStudy (anonymous):

Now\[{x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} \] we want to solve for x using the quadratic formula

OpenStudy (anonymous):

idk, it's going to take a while....

OpenStudy (anonymous):

It should take you a while at first... this really hard math but as long as you keep at it and solve for x you will get it

OpenStudy (jamesj):

Well, this is why you have these problems: to practice. We all have to do lots of these to get fast at them. So off you go!

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

Remember do it two times \[{x = \frac{{ - b- \sqrt {b^2 - 4ac} }}{{2a}}}\] \[{x = \frac{{ - b + \sqrt {b^2 - 4ac} }}{{2a}}}\]

OpenStudy (anonymous):

where does A=-2 go? on the outside right?

OpenStudy (anonymous):

nm I found it

OpenStudy (anonymous):

so it shud look like -3+- √3^2 4(-2)(-9)/2(-2)?

OpenStudy (anonymous):

???

OpenStudy (jamesj):

\[ \frac{-3 \pm \sqrt{3^2 - 4(-2)(-9)}}{2(-2)} \] yes. Now simplify.

OpenStudy (anonymous):

okay ty

OpenStudy (anonymous):

okay so now it's -3 +-√9 +72 /-4?

OpenStudy (jamesj):

sign error. Check your calculation.

OpenStudy (anonymous):

hint: it's inside the parentheses the error

OpenStudy (anonymous):

why what's wrong?

OpenStudy (anonymous):

okay anyway , now it's -3 +-√81/-4?

OpenStudy (anonymous):

\[-a*-b=+c\]

OpenStudy (jamesj):

Nope. --- = -, not --- = +

OpenStudy (anonymous):

isk what that means? can you expain?

OpenStudy (anonymous):

That means negative times a negative is a positive but inside the parentheses you have negative times negative times negative

OpenStudy (anonymous):

so two negatives will give you a positive times another negative will give you a negative

OpenStudy (anonymous):

so negative 72?

OpenStudy (anonymous):

\[{x = \frac{{ - 3 \pm \sqrt {3^2 - 4(-2)(-9)} }}{{2(-2)}}}\] \[{x = \frac{{ - 3 \pm \sqrt {9 - 72} }}{{-4}}}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ty.

OpenStudy (anonymous):

so now it'a -3+-√-63 /-4...now wat?

OpenStudy (anonymous):

it's*

OpenStudy (anonymous):

\[{x = \frac{{ - 3 \pm \sqrt {-63} }}{{-4}}}\]

OpenStudy (anonymous):

From there we want to get the complex number

OpenStudy (anonymous):

We get it from the negative sign inside the square root

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

what does that mean andf where did that come from?

OpenStudy (anonymous):

\[x=\frac{-3 \pm \sqrt{-63}}{-4}\] \[x=\frac{-3 \pm i \sqrt{63}}{-4}\]

OpenStudy (anonymous):

okay now what do we square root of 63?

OpenStudy (anonymous):

I really don't know what it means or how to explain it that well but basically the \[\sqrt{-1}=i\] and yeah you are right we can square root 63 now because now we don't have to worry about the negative value

OpenStudy (anonymous):

If you want to learn more about complex numbers you just wikipedia it.

OpenStudy (anonymous):

okay , thank you for all ur help :D

OpenStudy (anonymous):

\[x=\frac{-3 \pm i 3\sqrt{7}}{-4}\] \[x=\frac{-3}{-4} \pm \frac{3i \sqrt{7}}{-4} \] \[x=\frac{3}{4} \pm \frac{3i \sqrt{7}}{-4} \] \[x=\frac{3}{4} + \frac{3i \sqrt{7}}{-4} and, x=\frac{3}{4} - \frac{3i \sqrt{7}}{-4}\]

OpenStudy (anonymous):

that's noot one of my choices?

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

The problem is that we might of simplified it differently. I mean 3/4 actually equals 0.75 and you can also change the squareroot of 7 so that might be it

OpenStudy (anonymous):

x=3+-i√29/2 < and then the same but just -3 and then x=3+-i√23 and then the same just with -3 thses are my choies